OpenStudy (mpari13):

quick clarification: NO MATH! I feel like im getting confused can someone explain, when the radius value is decreased, how does the centripetal force change?

OpenStudy (anonymous):

\[F_{Centripetal} = \frac{mv_{T}^2}{r}\] Where Vt is the tangential velocity. So, if m and v stay the same, but we decrease r by half, what happens to F?

OpenStudy (mpari13):

it would get larger right?

OpenStudy (anonymous):

I know you said no math, but this is the clearest way that I know to show you. Let's say that we were a distance a from the pivot point. That is, r=a. Our centripetal force would be: \[F = \frac{mv^2}{a}\] Easy enough, right? Let's see what happens when we decrease this distance by half. r = (1/2)a \[F = \frac{mv^2}{\frac{1}{2}a} = \frac{2mv^2}{a}\] In general, we could summarize the behavior of F as we change r by just "ignoring" the things that aren't changing: \[F \propto \frac{1}{r}\]or equivalently: \[F \propto r^{-1}\] where the "fish" sign means "is proportional to" or "varies as" So F is proportional to one over r. So now, if we want to see what will happen to F if we decrease (or increase) r by some factor, we just throw it in for r. Let's look at one third as an example. We want to decrease r to one third of it's current amount. What will happen to F? \[F \propto (\frac{1}{3})^{-1} = 3\] So F will be increased by a factor of 3.

OpenStudy (anonymous):

Make sense?

OpenStudy (mpari13):

Makes sense, the Fc and radius are inversely proportional, so as radius decrease force increases.